# Fixing It in Post: A Comparative Study of LLM Post-Training Data Quality and Model Performance

### v2

@@ 1 Introduction
As large language models (LLMs) models continue to grow in complexity, so do their training requirements, necessitating ever-larger datasets with each new model iteration [1, 2, 3]. While pretraining LLMs on large, general‐purpose corpora is now well understood [3, 4, 2, 5, 1, 6], recent work has shifted toward post‐training, which includes supervised fine-tuning (SFT), reinforcement learning (RL), and task-specific fine-tuning such as domain adaptation [7, 8, 9, 10].
Carefully curated post-training datasets play a critical role in ensuring high downstream task performance, instruction following, and advanced reasoning. Nevertheless, the majority of post-training corpora remain proprietary, restricted by commercial licensing or intellectual‐property concerns, and are therefore unavailable for public scrutiny and reuse. This has motivated state-of-the-art research on synthetic data generation, the development of open-source large-scale post-training datasets, and the design of effective post-training recipes [11, 12, 13, 14, 7, 15, 16].
@@
Another challenge with these datasets is the lack of transparent documentation regarding their curation processes. General steps are typically addressed briefly and critical details, particularly those concerning the creation of dataset mixtures, are often vaguely described. While recent works [7, 13] take giant strides in enhancing the transparency of post-training datasets and recipes, obscurity still remains for several crucial aspects. For instance, there is often a lack of details on which exact ablations were conducted to design mixture ratios, rendering the approach neither fully replicable nor sufficiently insightful to guide future dataset curation efforts.
Thus, in this paper, we adopt a principled and openly insightful approach to developing and evaluating post-training datasets, an effort that, to the best of our knowledge, is the first of its kind. To this end, we focus on two of the largest openly available SFT mixtures [^1] While post-training typically includes both SFT and RL, we restrict our attention to SFT in this paper. See App. A for more details. from recent works: Tulu-3-SFT-Mix [7] (referred to as Tulu) and SmolTalk [9].
@@
  • Performance Evaluation: We conduct the first side-by-side comparison of recent open-source SFT data mixtures: Tulu [7], SmolTalk [13], and Orca-AgentInstruct [16]. By fixing the model architecture and training hyperparameters, we enable a clean comparison of dataset performance across 14 LLM benchmarks, including those from popular Open LLM Leaderboards [17, 18]. We identify key differences and performance gaps for specialized skills such as coding, math, and instruction following, illuminating strengths and weaknesses of each mixture.
  • Quality Annotations: To systematically drive data mixture decisions, we require detailed and standardized annotations of samples, a practice currently uncommon among the open-source post-training community. To this end, we leverage the Magpie framework [14] and annotate each Tulu and SmolTalk data sample along multiple dimensions, including conversational structure, prompt and response quality, and task categorization. These annotations provide concrete insights into dataset composition and support informed decision making for performant data mixtures.
@@ 2 Background and Motivation
To the best of our knowledge, a direct side-by-side dissection of two flagship post-training datasets such as Tulu and SmolTalk has not been previously conducted, primarily due to significant compute requirements, particularly for large-scale SFT and extensive data annotations. We specifically select Tulu and SmolTalk (see App. B for more details on dataset compositions) due to their widespread adoption and the strong empirical performance demonstrated by their respective post-trained LLMs in recent benchmarks [7, 9].
@@
Tulu. Lambert et al. [7] designed Tulu to advance broad-spectrum reasoning in medium-sized Llama models. They begin by filtering existing instruction corpora for (i) diverse real-user requests (e.g., WildChat [8], OpenAssistant [19]), and (ii) specialized skills (e.g., OpenMath-Instruct [10]). Residual gaps in instruction following, math, coding, and safety are filled with GPT-4o [20] generated prompts produced through persona-based prompting [21]. After n-gram decontamination and heuristic quality filtering, the final release contains 0.94 million high-quality pairs spanning seven broad domains (knowledge, math, reasoning, coding, safety, instruction following, and multilingual).
@@
SmolTalk. Allal et al. [9] pursue a complementary objective to Tulu by building small models that deliver rich, multi-turn conversations without requiring large compute budgets. Their SFT mixture therefore focuses on conversational depth and pragmatic rewriting, and comprises roughly 1.04 million examples. In their curation process, they first mix Magpie-Ultra [14] which provides high-quality multi-turn prompts generated via an enhanced two-step Magpie procedure on a stronger teacher model. Second, three synthetic task-oriented subsets (Smol-Constraint, Smol-Summarization, and Smol-Rewrite [9]) are produced with targeted system prompts on Qwen2.5-72B-Instruct [22]. Third, equation-heavy math corpora [23, 24] are added. Finally, they add code-alignment and long-context resources [25, 26, 12] to the mixture. All subsets undergo a similar deduplication, quality filtering, and n-gram decontamination as in Tulu.
@@
In addition to Tulu and SmolTalk, we also acknowledge earlier, similar-sized influential post-training mixtures such as Orca [15], which similarly covers tasks ranging from creative writing and text editing to coding and reading comprehension. However, initial results indicate that the more recent Tulu and SmolTalk datasets consistently outperform Orca across all evaluated tasks. Thus, we omit further investigation into Orca.
@@
We present the corresponding SFT results in Table 1, where we fine-tune Llama-3.1-8B [27] and SmolLM2-1.7B [9] on Tulu, SmolTalk, and Orca, respectively, and evaluate both models on popular OpenLLM Leaderboard benchmarks (see App. E for the detailed fine-tuning and evaluation setup). We select these two models because they were used by Lambert et al. [7] to train the Llama-3.1-Tulu-3-8B model on the Tulu dataset and by Allal et al. [9] to train the SmolLM2-1.7B-Instruct model on SmolTalk. This choice allows us to validate our pipelines and ensure parity with prior work.
@@
We evaluate performance across 12 tasks drawn from Open LLM Leaderboard V1 and Leaderboard V2, as well as two code generation tasks (HumanEval and HumanEval+), using the LM Evaluation Harness framework [28]. We report the average scores for Leaderboard V1 and Leaderboard V2 along with the overall average score across all 14 benchmarks. For Llama, SFT on SmolTalk outperforms Tulu on both LLM leaderboard benchmarks, however, falls behind in code benchmarks, where Tulu slightly pulls ahead on both benchmarks. Overall, fine-tuning with either dataset yields substantial improvements compared to the baseline. For SmolLM, both SmolTalk and Tulu achieve similar performance on the OpenLLM Leaderboard benchmarks. However, performance is equally low on both code benchmarks, reflecting SmolLM’s smaller size and its design focus on conversational rather than coding tasks. As noted earlier, both Tulu and SmolTalk consistently outperform Orca across all benchmarks, reflecting stronger data curation and higher corresponding task coverage.
@@ 3.1 Unified Magpie Annotations
To enable direct comparability between Tulu and SmolTalk, we annotate (tag) every data sample using Magpie with Llama-3.3-70B-Instruct [27] as the judge model (see App. C.1). We find that Llama-generated annotations are reliable and align with human judgment (see App. C.1.3 and App. C.1.4). Magpie annotates each sample with structured tags for Task Category (12 classes), Input Quality (rated from very poor to excellent), Response Quality (termed Instruct Reward, rated from 0 to 5 for multi-turn and as a real number for single-turn), Safety (assessed via Llama-Guard 2 [29]), Language, and query Difficulty. We further extend Magpie’s original annotation set by explicitly capturing the conversation structure (single-turn vs. multi-turn) and retain important metadata from the original datasets (e.g., unique sample identifiers), making our annotated versions reusable for future research.
@@ 4 Leveraging Annotations to Design Data Curation Recipes
Ablation Setup. We evaluate our curation recipes through ablation experiments. We use stratified sampling to extract a representative subsample of approximately 10% (about 100k examples) from each of the original Tulu and SmolTalk datasets, resulting in subsets Tulu-100k and SmolTalk-100k. A subsample size of 10% is chosen to make training more computationally efficient while preserving performance trends, as similarly demonstrated in the original Tulu experiments [7]. We apply our curation criteria to these subsamples, selecting high-quality instances to form a new data mixture.
@@ 4.1 Quality-Based Curation Recipe
Diversity Analysis. In our preliminary analysis, instruction following emerged as a critical capability influencing performance on other benchmarks (similar observations are also reported in [30, 7]). By filtering the annotated Tulu and SmolTalk datasets for sources explicitly containing instruction following tasks, we observe that many such examples fall into the categories advice seeking, information seeking, creative writing, and reasoning. Fig. 6 illustrates the resulting task diversity across the considered datasets, highlighting significant reductions in these instruction-rich categories within TuluTalk-80k. Notably, the proportion of information seeking samples drops to 12% compared to 20% in SmolTalk and 25% in Tulu. This confirms that our quality-based curation recipe requires additional task-aware adaptation to include more instruction following examples.
@@ 5 Results on Full Datasets and Discussion
Furthermore, to assess generalizability across model scales and architectures, we conduct additional experiments using Qwen2.5-0.5B and Qwen2.5-3B [22], as well as SmolLM3-3B [31]. Table 4 reports the Open LLM Leaderboard and overall average scores across all benchmarks for each model. The results show that TuluTalk consistently outperforms Tulu and SmolTalk across all models, confirming our prior analysis and demonstrating robust cross-model generalization. Comprehensive evaluations for each model are provided in App. E.3.4. In addition, App. E.3.5 presents results for the Llama model fine-tuned with Direct Preference Optimization (DPO) [32], further demonstrating that the performance gains observed under SFT carry over to the DPO setting.
@@ A.1 Post-Training Workflow
The goal of preference fine-tuning is is to align the model’s output distribution with human preferences or task-specific objectives. This is typically done by guiding the model using a reward model or preference signal to prefer helpful, harmless, and honest completions. Popular algorithms for preference tuning include Proximal Policy Optimization (PPO) [33], Group Relative Policy Optimization (GRPO) [34], and Direct Preference Optimization (DPO) [32].
@@
Recent work has explored reinforcement learning and preference-based methods to enhance deep thinking capabilities in LLMs, such as multi-hop reasoning [35], chain-of-thought generation [36], tool use [37], and debate-style deliberation [38]. These methods typically rely on reward models or heuristic scoring to reward structured reasoning behavior that extends beyond surface-level fluency. Corresponding reasoning-centric datasets have emerged as well [39, 40, 41, 42, 43], which introduce task formats that elicit step-by-step thought processes.
@@ A.2 Focus on SFT
The primary goal of this paper is to analyze the quality and composition of training datasets while keeping the training procedure fixed. In particular, we focus on SFT because the performance of SFT-tuned models is largely governed by the structure and quality of the data mixture rather than by training algorithmic nuances. Furthermore, most open-source SFT pipelines follow similar training setups, whereas preference fine-tuning introduces additional complexity: the training algorithm (e.g., PPO, DPO, RLVR) directly determines the type and structure of data it can effectively utilize. For example, PPO requires preference pairs to train a reward model, followed by policy rollouts for fine-tuning [33]. DPO, by contrast, directly trains on preference pairs without requiring policy rollouts or a reward model [32]. Other methods like Reinforcement Learning with Verifiable Rewards (RLVR) require examples with verifiable numeric rewards [7].
The diversity of preference tuning recipes and their data format dependencies thus makes clean cross-method comparisons challenging. Indeed, designing and evaluating preference-based training pipelines is itself an active research area [7, 44, 45]. We leave the analysis of data quality under different alignment strategies to future work (see App. F).
@@ A.3 Related SFT Datasets
Tulu [7] and SmolTalk [9], investigated in this paper, are two of the most recent and widely used open-source datasets for SFT post-training of LLMs. We focus on these two datasets due to their strong reported performance over prior SFT datasets across a broad range of benchmarks when used to train the respective models introduced in their original papers.
Several other SFT datasets have been proposed in recent years, including Orca [16], OpenHermes [12], LongAlign [46], UltraFeedback [47], MAmmoTH2 [48], DaringAntEater [49] Magpie-Pro [14] RLHFlow-SFT-V2 [50]. While many of these datasets provide valuable capabilities, such as long-context support, synthetic feedback signals, or broad coverage across domains, Tulu and SmolTalk remain highly competitive, achieving significantly stronger performance across instruction following, reasoning, and code tasks [7, 9].
@@
Furthermore, as shown in App. B, both Tulu and SmolTalk include carefully curated subsets drawn from several of the datasets mentioned above, particularly from OpenHermes2.5 [12], Smol-Magpie-Ultra [9], OpenAssistant [19], and UltraFeedback [47].

### v3

@@ 1 Introduction
As large language models (LLMs) models continue to grow in complexity, so do their training requirements, necessitating ever-larger datasets with each new model iteration [li2024datacomp, penedo2024fineweb, weber2024redpajama]. While pretraining LLMs on large, general‐purpose corpora is now well understood [weber2024redpajama, penedo2023refinedweb, penedo2024fineweb, soldaini2024dolma, li2024datacomp, txt360data2024], recent work has shifted toward post‐training, which includes supervised fine-tuning (SFT), reinforcement learning (RL), and task-specific fine-tuning such as domain adaptation [lambert2025tulu3pushingfrontiers, zhao2024wildchat, allal2025smollm2smolgoesbig, toshniwal2024openmathinstruct2acceleratingaimath].
Carefully curated post-training datasets play a critical role in ensuring high downstream task performance, instruction following, and advanced reasoning. Nevertheless, the majority of post-training corpora remain proprietary, restricted by commercial licensing or intellectual‐property concerns, and are therefore unavailable for public scrutiny and reuse. This has motivated state-of-the-art research on synthetic data generation, the development of open-source large-scale post-training datasets, and the design of effective post-training recipes [benallal2024cosmopedia, OpenHermes2.5, allal2024SmolLM, xu2024magpie, lambert2025tulu3pushingfrontiers, mukherjee2023orca, mitra2024agentinstruct].
@@
Another challenge with these datasets is the lack of transparent documentation regarding their curation processes. General steps are typically addressed briefly and critical details, particularly those concerning the creation of dataset mixtures, are often vaguely described. While recent works [lambert2025tulu3pushingfrontiers, allal2024SmolLM] take giant strides in enhancing the transparency of post-training datasets and recipes, obscurity still remains for several crucial aspects. For instance, there is often a lack of details on which exact ablations were conducted to design mixture ratios, rendering the approach neither fully replicable nor sufficiently insightful to guide future dataset curation efforts.
Thus, in this paper, we adopt a principled and openly insightful approach to developing and evaluating post-training datasets, an effort that, to the best of our knowledge, is the first of its kind. To this end, we focus on two of the largest openly available SFT mixtures [^1] While post-training typically includes both SFT and RL, we restrict our attention to SFT in this paper. See App. A for more details. from recent works: Tulu-3-SFT-Mix [lambert2025tulu3pushingfrontiers] (referred to as Tulu) and SmolTalk [allal2025smollm2smolgoesbig].
@@
  • Performance Evaluation: We conduct the first side-by-side comparison of recent open-source SFT data mixtures: Tulu [lambert2025tulu3pushingfrontiers], SmolTalk [allal2024SmolLM], and Orca-AgentInstruct [mitra2024agentinstruct]. By fixing the model architecture and training hyperparameters, we enable a clean comparison of dataset performance across 14 LLM benchmarks, including those from popular Open LLM Leaderboards [leaderboardv1, open-llm-leaderboard-v2]. We identify key differences and performance gaps for specialized skills such as coding, math, and instruction following, illuminating strengths and weaknesses of each mixture.
  • Quality Annotations: To systematically drive data mixture decisions, we require detailed and standardized annotations of samples, a practice currently uncommon among the open-source post-training community. To this end, we leverage the Magpie framework [xu2024magpie] and annotate each Tulu and SmolTalk data sample along multiple dimensions, including conversational structure, prompt and response quality, and task categorization. These annotations provide concrete insights into dataset composition and support informed decision making for performant data mixtures.
@@ 2 Background and Motivation
To the best of our knowledge, a direct side-by-side dissection of two flagship post-training datasets such as Tulu and SmolTalk has not been previously conducted, primarily due to significant compute requirements, particularly for large-scale SFT and extensive data annotations. We specifically select Tulu and SmolTalk (see App. B for more details on dataset compositions) due to their widespread adoption and the strong empirical performance demonstrated by their respective post-trained LLMs in recent benchmarks [lambert2025tulu3pushingfrontiers, allal2025smollm2smolgoesbig].
@@
Tulu. lambert2025tulu3pushingfrontiers designed Tulu to advance broad-spectrum reasoning in medium-sized Llama models. They begin by filtering existing instruction corpora for (i) diverse real-user requests (e.g., WildChat [zhao2024wildchat], OpenAssistant [kopf2023openassistant]), and (ii) specialized skills (e.g., OpenMath-Instruct [toshniwal2024openmathinstruct2acceleratingaimath]). Residual gaps in instruction following, math, coding, and safety are filled with GPT-4o [openai2024gpt4ocard] generated prompts produced through persona-based prompting [ge2024scaling]. After n-gram decontamination and heuristic quality filtering, the final release contains 0.94 million high-quality pairs spanning seven broad domains (knowledge, math, reasoning, coding, safety, instruction following, and multilingual).
@@
SmolTalk. allal2025smollm2smolgoesbig pursue a complementary objective to Tulu by building small models that deliver rich, multi-turn conversations without requiring large compute budgets. Their SFT mixture therefore focuses on conversational depth and pragmatic rewriting, and comprises roughly 1.04 million examples. In their curation process, they first mix Magpie-Ultra [xu2024magpie] which provides high-quality multi-turn prompts generated via an enhanced two-step Magpie procedure on a stronger teacher model. Second, three synthetic task-oriented subsets (Smol-Constraint, Smol-Summarization, and Smol-Rewrite [allal2025smollm2smolgoesbig]) are produced with targeted system prompts on Qwen2.5-72B-Instruct [yang2024qwen2]. Third, equation-heavy math corpora [numina_math_datasets, yu2023metamath] are added. Finally, they add code-alignment and long-context resources [wei2024selfcodealign, liu2024apigen, OpenHermes2.5] to the mixture. All subsets undergo a similar deduplication, quality filtering, and n-gram decontamination as in Tulu.
@@
In addition to Tulu and SmolTalk, we also acknowledge earlier, similar-sized influential post-training mixtures such as Orca [mukherjee2023orca], which similarly covers tasks ranging from creative writing and text editing to coding and reading comprehension. However, initial results indicate that the more recent Tulu and SmolTalk datasets consistently outperform Orca across all evaluated tasks. Thus, we omit further investigation into Orca.
@@
We present the corresponding SFT results in Table 1, where we fine-tune Llama-3.1-8B [grattafiori2024llama3herdmodels] and SmolLM2-1.7B [allal2025smollm2smolgoesbig] on Tulu, SmolTalk, and Orca, respectively, and evaluate both models on popular OpenLLM Leaderboard benchmarks (see App. E for the detailed fine-tuning and evaluation setup). We select these two models because they were used by lambert2025tulu3pushingfrontiers to train the Llama-3.1-Tulu-3-8B model on the Tulu dataset and by allal2025smollm2smolgoesbig to train the SmolLM2-1.7B-Instruct model on SmolTalk. This choice allows us to validate our pipelines and ensure parity with prior work.
@@
We evaluate performance across 12 tasks drawn from Open LLM Leaderboard V1 and Leaderboard V2, as well as two code generation tasks (HumanEval and HumanEval+), using the LM Evaluation Harness framework [lm-eval-harness]. We report the average scores for Leaderboard V1 and Leaderboard V2 along with the overall average score across all 14 benchmarks. For Llama, SFT on SmolTalk outperforms Tulu on both LLM leaderboard benchmarks, however, falls behind in code benchmarks, where Tulu slightly pulls ahead on both benchmarks. Overall, fine-tuning with either dataset yields substantial improvements compared to the baseline. For SmolLM, both SmolTalk and Tulu achieve similar performance on the OpenLLM Leaderboard benchmarks. However, performance is equally low on both code benchmarks, reflecting SmolLM’s smaller size and its design focus on conversational rather than coding tasks. As noted earlier, both Tulu and SmolTalk consistently outperform Orca across all benchmarks, reflecting stronger data curation and higher corresponding task coverage.
@@ 3.1 Unified Magpie Annotations
To enable direct comparability between Tulu and SmolTalk, we annotate (tag) every data sample using Magpie with Llama-3.3-70B-Instruct [grattafiori2024llama3herdmodels] as the judge model (see App. C.1). We find that Llama-generated annotations are reliable and align with human judgment (see App. C.1.3 and App. C.1.4). Magpie annotates each sample with structured tags for Task Category (12 classes), Input Quality (rated from very poor to excellent), Response Quality (termed Instruct Reward, rated from 0 to 5 for multi-turn and as a real number for single-turn), Safety (assessed via Llama-Guard 2 [metallamaguard2]), Language, and query Difficulty. We further extend Magpie’s original annotation set by explicitly capturing the conversation structure (single-turn vs. multi-turn) and retain important metadata from the original datasets (e.g., unique sample identifiers), making our annotated versions reusable for future research.
@@ 4 Leveraging Annotations to Design Data Curation Recipes
Ablation Setup. We evaluate our curation recipes through ablation experiments. We use stratified sampling to extract a representative subsample of approximately 10% (about 100k examples) from each of the original Tulu and SmolTalk datasets, resulting in subsets Tulu-100k and SmolTalk-100k. A subsample size of 10% is chosen to make training more computationally efficient while preserving performance trends, as similarly demonstrated in the original Tulu experiments [lambert2025tulu3pushingfrontiers]. We apply our curation criteria to these subsamples, selecting high-quality instances to form a new data mixture.
@@ 4.1 Quality-Based Curation Recipe
Diversity Analysis. In our preliminary analysis, instruction following emerged as a critical capability influencing performance on other benchmarks (similar observations are also reported in [cohere2025commanda, lambert2025tulu3pushingfrontiers]). By filtering the annotated Tulu and SmolTalk datasets for sources explicitly containing instruction following tasks, we observe that many such examples fall into the categories advice seeking, information seeking, creative writing, and reasoning. Fig. 6 illustrates the resulting task diversity across the considered datasets, highlighting significant reductions in these instruction-rich categories within TuluTalk-80k. Notably, the proportion of information seeking samples drops to 12% compared to 20% in SmolTalk and 25% in Tulu. This confirms that our quality-based curation recipe requires additional task-aware adaptation to include more instruction following examples.
@@ 5 Results on Full Datasets and Discussion
Furthermore, to assess generalizability across model scales and architectures, we conduct additional experiments using Qwen2.5-0.5B and Qwen2.5-3B [yang2024qwen2], as well as SmolLM3-3B [bakouch2025smollm3]. Table 4 reports the Open LLM Leaderboard and overall average scores across all benchmarks for each model. The results show that TuluTalk consistently outperforms Tulu and SmolTalk across all models, confirming our prior analysis and demonstrating robust cross-model generalization. Comprehensive evaluations for each model are provided in App. E.3.4. In addition, App. E.3.5 presents results for the Llama model fine-tuned with Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage], further demonstrating that the performance gains observed under SFT carry over to the DPO setting.
@@ A.1 Post-Training Workflow
The goal of preference fine-tuning is is to align the model’s output distribution with human preferences or task-specific objectives. This is typically done by guiding the model using a reward model or preference signal to prefer helpful, harmless, and honest completions. Popular algorithms for preference tuning include Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms], Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], and Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage].
@@
Recent work has explored reinforcement learning and preference-based methods to enhance deep thinking capabilities in LLMs, such as multi-hop reasoning [yang2024largelanguagemodelslatently], chain-of-thought generation [hao2024traininglargelanguagemodels], tool use [qin2023toolllmfacilitatinglargelanguage], and debate-style deliberation [du2023improvingfactualityreasoninglanguage]. These methods typically rely on reward models or heuristic scoring to reward structured reasoning behavior that extends beyond surface-level fluency. Corresponding reasoning-centric datasets have emerged as well [bercovich2025llamanemotronefficientreasoningmodels, sky_t1_2025, bespoke_stratos, slam-distillation-from-r1, still], which introduce task formats that elicit step-by-step thought processes.
@@ A.2 Focus on SFT
The primary goal of this paper is to analyze the quality and composition of training datasets while keeping the training procedure fixed. In particular, we focus on SFT because the performance of SFT-tuned models is largely governed by the structure and quality of the data mixture rather than by training algorithmic nuances. Furthermore, most open-source SFT pipelines follow similar training setups, whereas preference fine-tuning introduces additional complexity: the training algorithm (e.g., PPO, DPO, RLVR) directly determines the type and structure of data it can effectively utilize. For example, PPO requires preference pairs to train a reward model, followed by policy rollouts for fine-tuning [schulman2017proximalpolicyoptimizationalgorithms]. DPO, by contrast, directly trains on preference pairs without requiring policy rollouts or a reward model [rafailov2024directpreferenceoptimizationlanguage]. Other methods like Reinforcement Learning with Verifiable Rewards (RLVR) require examples with verifiable numeric rewards [lambert2025tulu3pushingfrontiers].
The diversity of preference tuning recipes and their data format dependencies thus makes clean cross-method comparisons challenging. Indeed, designing and evaluating preference-based training pipelines is itself an active research area [lambert2025tulu3pushingfrontiers, ivison2024unpackingdpoppo, winata2024preferencetuning]. We leave the analysis of data quality under different alignment strategies to future work (see App. F).
@@ A.3 Related SFT Datasets
Tulu [lambert2025tulu3pushingfrontiers] and SmolTalk [allal2025smollm2smolgoesbig], investigated in this paper, are two of the most recent and widely used open-source datasets for SFT post-training of LLMs. We focus on these two datasets due to their strong reported performance over prior SFT datasets across a broad range of benchmarks when used to train the respective models introduced in their original papers.
Several other SFT datasets have been proposed in recent years, including Orca [mitra2024agentinstruct], OpenHermes [OpenHermes2.5], LongAlign [bai2024longalign], UltraFeedback [cui2024ultrafeedbackboostinglanguagemodels], MAmmoTH2 [yue2024mammoth2scaling], DaringAntEater [wang2024helpsteer2opensource] Magpie-Pro [xu2024magpie] RLHFlow-SFT-V2 [RLHFlow]. While many of these datasets provide valuable capabilities, such as long-context support, synthetic feedback signals, or broad coverage across domains, Tulu and SmolTalk remain highly competitive, achieving significantly stronger performance across instruction following, reasoning, and code tasks [lambert2025tulu3pushingfrontiers, allal2025smollm2smolgoesbig].
@@
Furthermore, as shown in App. B, both Tulu and SmolTalk include carefully curated subsets drawn from several of the datasets mentioned above, particularly from OpenHermes2.5 [OpenHermes2.5], Smol-Magpie-Ultra [allal2025smollm2smolgoesbig], OpenAssistant [kopf2023openassistant], and UltraFeedback [cui2024ultrafeedbackboostinglanguagemodels].